OpenStudy (anonymous):

Evaluate (f o g)(-12)

OpenStudy (anonymous):

\[f(x)=1-4x ^{2}, g(x)=\sqrt{1-2x}\]

OpenStudy (anonymous):

i don't get the (-12) part

OpenStudy (anonymous):

f o g = f(g(x)) = f( sqrt[1-2x] ) = 1 - 4*[sqrt{1-2x}] ^2 u will get an equation put -12 in that equation

OpenStudy (anonymous):

f o g means f o g(x) = ..... then f o g(-12) = put -12 instead of x in the resulting expresion

OpenStudy (anonymous):

so i'll just substitute x with -12

OpenStudy (anonymous):

ya but when u simplify the equation f o g (x) = f(g(x)) = f( sqrt[1-2x] ) = 1 - 4*[sqrt{1-2x}] ^2=1-4[1-2x] = 8x - 3 now put -12 so u get 8(-12) -3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

most welcome